OpenStudy (anonymous):

factor 27x^4-12 3^rad(x)-36/3^radx it's hard to communicate but the 3^rad(x) means x cubed :/

OpenStudy (anonymous):

So is it supposed to look like this?

OpenStudy (anonymous):

\[27x ^{4}-12\sqrt[3]{x}-\frac{ 36 }{ \sqrt[3]{x} }\]

OpenStudy (anonymous):

its supposed to look like 27x\[27x^{4}-12\sqrt[3]{x}-36/\sqrt[3]{x}\]